### v1

# Marathon: A Race Through the Realm of Long Context with Large Language Models

Abstract
Although there are currently many benchmarks available for evaluating the long context understanding and reasoning capability of large language models, with the expansion of the context window in these models, the existing long context benchmarks are no longer sufficient for evaluating the long context understanding and reasoning capability of large language models. In this paper, we have developed a fresh long context evaluation benchmark, which we name it Marathon in the form of multiple choice questions, inspired by benchmarks such as MMLU, for assessing the long context comprehension capability of large language models quickly, accurately, and objectively. We have evaluated several of the latest and most popular large language models, as well as three recent and effective long context optimization methods, on our benchmark. This showcases the long context reasoning and comprehension capabilities of these large language models and validates the effectiveness of these optimization methods. Marathon is available at https://huggingface.co/datasets/Lemoncoke/Marathon. [^1] We also provide an online evaluation website: https://openbenchmark.online/marathon
1 Introduction
Figure 1: An example of a test case in benchmark. The context is truncated for display purposes.
With the rapid development of artificial intelligence technology, large language models represented by GPT-4 (OpenAI, 2023b) have demonstrated powerful capabilities. In addition to the well-known ChatGPT, the application of large language models has gradually extended to various industries. The large model in the medical field, MedGPT, can help people consult medical-related questions, while the large model in the legal field, LawGPT (Nguyen, 2023), can assist people in legal consultations, significantly reducing the barriers and costs for people to obtain knowledge. The large model in the programming field, Github Copilot (Chen et al., 2021), can help developers complete their work faster, greatly improving productivity.
However, current large language models can only provide answers to users’ questions and requests within a relatively small context window. There are two main reasons for this limitation. Firstly, large language models still lack the ability to extract crucial information from long contexts. As described in Liu et al. (2023) large language models can only effectively retrieve key information when it is located at the beginning or end of the prompt. When the key information is in the middle of the prompt, it becomes difficult to retrieve successfully. This phenomenon becomes more evident as the context length increases. Secondly, longer contexts require more computational resources and time for large language models to provide feedback, which not only increases the hardware requirements but also reduces the user experience. Therefore, the ability of large language models to handle long contexts becomes increasingly important, and comprehensive evaluation of their performance in this regard is crucial. Existing long context benchmarks, such as Longbench (Bai et al., 2023b), still employ the F1 score as the evaluation metric, which measures the similarity between the generated response of the large language model and the candidate answers. However, the generated content of large language models is highly diverse, and the candidate answers cannot cover all possible correct responses, making the evaluation of large language models less accurate.
To address these challenges, we propose a new long context benchmark called Marathon based on Loogle (Li et al., 2023b), which includes 6 evaluation tasks for assessing the ability of large language models. The context lengths in the benchmark range from 60k to 260k+. For each long context, there is a related question and four carefully crafted answer options. The large language model needs to provide the correct answer option based on the given long context.
We conducted evaluations on Marathon using various large language models with different parameter sizes and context windows, combined with different methods of compressing long contexts. We found that the long context processing capabilities of most large language models still need improvement. In fact, the 7B Mistral (Jiang et al., 2023a) model outperformed the 70B Beluga (Mahan et al., 2023) model in terms of long context understanding. Some compression methods for long contexts were found to be effective in improving large language models’ understanding of long contexts in certain cases.
2 Related Work
2.1 Flash Attention
The training and inference of ultra-long sequences have been hindering the further development of large language models. In traditional Attention mechanisms, the GPU memory consumption of Attention Scores is proportional to the square of the sequence length. Therefore, when the sequence length reaches a certain threshold, both training and inference will encounter out-of-memory (OOM) issues. The emergence of Flash Attention (Dao et al., 2022) allows us to avoid computing the Attention Scores matrix and directly compute the final Attention Output result, greatly reducing the GPU memory required during model training and inference. Moreover, thanks to the reduction in the number of memory accesses between HMB (high-bandwidth memory) and GPU SRAM (static random-access memory) brought by Flash Attention, the speed of model training and inference is accelerated. Flash Attention 2 (Dao, 2023) further optimizes the implementation of Flash Attention, further accelerating the speed of model computation. Currently, Flash Attention 2 (Dao, 2023) has become a key technology for optimizing the training and inference process of large language models, playing a crucial role in modeling and inferring ultra-long sequences.
2.2 Long Context Models
The ability of large language models for handling long contexts has become increasingly important. ChatGPT (OpenAI, 2023a) supports a window size of 16k, while GPT-4 supports a window size of 128k, and Claude-2.1 supports a window size of 200k [^2] https://www.anthropic.com/index/claude-2-1. Many open-source large language models have started to expand the size of their context window. Longchat (Li et al., 2023a) and MPT (Team, 2023) have achieved a window size of 16k, while Mistral (Jiang et al., 2023a) and Zephyr (Tunstall et al., 2023) have achieved a window size of 32k. By utilizing an adapted Rotary Embedding (Su et al., 2022) and sliding window (Beltagy et al., 2020) during fine-tuning, MistralLite, based on Mistral, has achieved a window size of 128k, enabling large language models to handle even longer contextual information.
2.3 Prompt Compression
Although larger context windows enable large language models to handle longer contextual information, processing long-context information requires a significant amount of computing resources and places high demands on hardware. It also necessitates longer computational time, even in the inference stage. Therefore, some methods like LLMLingua (Jiang et al., 2023c) and LongLLMLingua (Jiang et al., 2023b) have been proposed to compress long contexts.
2.4 Retrieval Augmented Generation
Retrieval Augmented Generation (RAG) was originally proposed and applied to NLP tasks in (Lewis et al., 2020), and it has now become a mainstream method for improving the generation capability of large language models. RAG can extract the most relevant data from external knowledge bases and hand it over to the large language model for processing. This can alleviate the hallucination problem of large language models and enable people to trace the source of the content generated by large language models, ensuring the reliability of the generated content. Additionally, RAG can also be used to extract information from long documents that is most relevant to the user’s query. This ensures that key information required to provide correct answers to questions is not lost while reducing the length of the context. Many projects such as longchain (longchain, 2022) and LlamaIndex (Liu, 2022) have achieved significant progress in combining RAG with large language models, greatly facilitating related research in this direction.
2.5 Long Context Benchmarks
There have been many recent benchmarks used to assess the long context processing ability of large language models, such as Loogle (Li et al., 2023b) and LongBench (Bai et al., 2023b). Liu et al. (2023) on the other hand, noticed that the position of key information in long contexts greatly affects the capability of large language models to correctly understand and process text. Therefore, they used the NaturalQA (Kwiatkowski et al., 2019) dataset to construct a new benchmark to test the impact of different positions of key information in long context on the text processing capability of large language models.
Although Loogle (Li et al., 2023b) and LongBench (Bai et al., 2023b) have constructed a relatively comprehensive set of evaluation tasks, the evaluation metrics used are still F1-score, Bleu or Rouge, which cannot accurately evaluate the ability of large language models to handle and understand long contexts.
Task No. Samples
Comprehension and Reasoning 357
Multiple Information Retreival 341
Timeline Reorder 152
Computation 97
Passage Retrieval 300
Short Dependency QA 283
Total 1530
Table 1: Statistics of Marathon.
Figure 2: The distribution of context length in Marathon.
3 Marathon
Currently, the mainstream benchmarks for evaluating large language models mostly adopt a multiple-choice question format, such as MMLU (Hendrycks et al., 2021) and C-Eval (Huang et al., 2023). The use of multiple-choice questions can effectively avoid scenarios where large language models generate correct answers but receive low scores due to the lack of relevant representations in the reference answers, or generate correct answers but receive high scores due to their similarity with the content of the reference answers. Therefore, based on Loogle (Li et al., 2023b) and Longbench (Bai et al., 2023b), we have constructed a multiple-choice long-text benchmark to more reasonably assess the long-context comprehension ability of large language models.
3.1 Overview
The Marathon benchmark consists of 6 tasks: Comprehension and Reasoning, Multiple Information Retrieval, Timeline Reorder, Computation, Passage Retrieval, and Short Dependency QA. These tasks can be categorized into four types based on the nature of the questions: Question Answering, Timeline Reorder, Computation, and Passage Retrieval. Figure 1 shows sample test cases for each type of question. Table 1 shows the quantity information of test samples for each task.
3.2 Construction
All the test samples in the benchmark are in the form of multiple-choice questions, with each question containing one correct answer option and several distractor options. We use GPT-4 to generate the distractor options for each question. For each question, we divide the long context into multiple fragments of length 12,000 and randomly select one fragment. We require GPT-4 to generate three distractor options based on the given context fragment, question, and correct answer. The purpose of this approach is to avoid using excessively long context that exceeds GPT-4’s context window, which may affect the accuracy of the generated results. By using shorter contexts, we can obtain distractor options that are more relevant to these shorter contexts. Finally, to ensure the effectiveness and accuracy of these distractor options, we manually verify the options of each test sample.
3.3 Question Answering
Comprehension and Reasoning, Multiple Information Retrieval and Short Dependency QA are all types of traditional question-answer formats. The difference lies in the fact that Comprehension and Reasoning, Multiple Information Retrieval are selected from the Long Dependency QA dataset in Loogle (Li et al., 2023b), while Short Dependency QA is selected from the Short Dependency QA dataset in Loogle (Li et al., 2023b). In the question answering tasks, each question is accompanied by a corresponding long context, and the large language model is required to infer the correct answer according to the long context. For the Short Dependency QA task, the relevant content for the correct answer is relatively concentrated within the long context. For Comprehension and Reasoning and Multiple Information Retrieval tasks, the content relevant to the correct answer is more scattered throughout the long context. Therefore, the large language model needs to possess strong long context understanding capability in order to solve the question correctly.
In the upper left of Figure 1, an example of a Question Answering task is provided. The question asks the large language model to answer a related question based on the content in the long context.
3.4 Timeline Reorder
Timeline Reorder task is a relatively novel question answering task. Unlike traditional question answering tasks, in the Timeline Reorder task, the question format requires large language models to sort a series of events described in a long context according to their chronological order. This task aims to examine the large language models’ understanding of temporal relationships. Due to the dispersed distribution of events that need to be sorted by chronological order in the long context, large language models not only need to possess a correct understanding of temporal order but also require strong long context processing capabilities to answer correctly, which makes it a challenging task.
In the upper right of Figure 1, an example of the Timeline Reorder task is provided. The question requires the large language model to sort three events mentioned in the long context according to their chronological order.
3.5 Computation
Computation task is also different from traditional question answering tasks. Its question format involves providing a question related to numerical computation and requires the large language model to perform numerical calculations based on relevant content in the long context. For example, it may require calculating the number of children a certain character has at a specific time point, considering that the long context describes the character’s life events, including the death of a child due to illness, which may affect the number of the character’s offspring at subsequent time points. Therefore, to answer this question correctly, the large language model not only needs to be able to perform ordinary numerical calculations but also needs to capture all the key information related to the question. Compared to traditional computation and question answering tasks, this task is more challenging and can better reflect the large language model’s capability to comprehend long context.
In the bottom left of Figure 1, an example of a Computation task is provided. The question requires the large language model to complete a numerical calculation question based on the content in the long context.
3.6 Passage Retrieval
Passage Retrieval task is one form of task in the Longbench (Bai et al., 2023b). In order to enhance the diversity of our benchmark tasks, we have sampled 300 test data from the Passage Retrieval task in Longbench (Bai et al., 2023b), and reformed them into multiple-choice format using the method mentioned above. We have incorporated this task into our benchmark. The Passage Retrieval task requires large language models to locate the paragraph in a long context that corresponds to the given description in the question. Since the test data of the Passage Retrieval task is sampled from Longbench (Bai et al., 2023b), there are some limitations in terms of context length and timeliness. However, it remains a highly valuable task format. In future work, we will update its content to make it more suitable for the current needs of evaluating large language models.
A sample of the Passage Retrieval task is provided in the bottom right of Figure 1. The task requires large language models to locate the paragraph in a long context that corresponds to the given description in the question.
Model Para. CW C&R MIR TR Com. PR SDQA Avg.
GPT-4 - 128K 77.03% 69.21% 69.08% 60.82% 100.00% 95.41% 78.59%
ChatGPT - 16K 62.18% 51.32% 19.74% 34.02% 95.67% 81.27% 57.37%
Vanilla
Chatglm 6B 32K 55.46% 46.63% 30.26% 37.11% 81.33% 79.51% 55.05%
Mistral 7B 32K 46.22% 41.94% 28.95% 23.71% 49.67% 48.41% 39.81%
Zephyr 7B 32K 41.46% 37.83% 33.24% 21.65% 47.67% 47.00% 37.97%
Longchat 13B 16K 37.25% 34.60% 28.29% 27.84% 42.00% 45.23% 35.87%
Qwen 14B 8K 45.38% 39.00% 26.32% 23.71% 56.33% 44.88% 39.27%
Yi 34B 200K 59.66% 47.21% 37.50% 36.08% 90.00% 65.02% 55.91%
Alfred 40B 8K 40.90% 39.30% 26.32% 20.62% 49.00% 47.70% 37.31%
Beluga 70B 4K 55.74% 43.70% 36.84% 36.08% 65.33% 59.36% 49.51%
Tulu2 70B 8K 46.50% 35.48% 30.26% 22.68% 46.33% 46.29% 37.92%
LongLLMLingua Compression
Chatglm 6B 32K 47.06% 37.54% 25.66% 22.68% 98.33% 56.18% 47.91%
Mistral 7B 32K 40.06% 31.38% 23.03% 27.84% 57.00% 42.76% 37.01%
Zephyr 7B 32K 30.81% 26.39% 23.68% 18.56% 54.00% 27.92% 30.23%
Longchat 13B 16K 37.82% 28.74% 26.32% 20.62% 61.67% 38.52% 35.61%
Qwen 14B 8K 42.58% 36.66% 27.63% 26.80% 88.67% 42.40% 44.12%
Yi 34B 200K 49.58% 42.23% 30.26% 22.68% 90.33% 56.89% 48.66%
Alfred 40B 8K 38.94% 32.84% 26.32% 29.90% 59.00% 45.94% 38.82%
Beluga 70B 4K 50.42% 42.82% 36.84% 27.84% 94.00% 63.60% 52.59%
Tulu2 70B 8K 45.94% 35.19% 34.87% 12.37% 98.00% 53.00% 46.56%
OpenAI Embedding Retrieval
Chatglm3 6B 32K 56.58% 43.40% 28.95% 28.87% 81.33% 66.78% 50.99%
Mistral 7B 32K 51.54% 47.21% 27.63% 27.84% 79.00% 67.84% 50.18%
Zephyr 7B 32K 52.38% 43.99% 28.29% 24.74% 76.67% 71.73% 49.63%
Longchat 13B 16K 38.10% 25.81% 19.08% 12.37% 43.00% 41.34% 29.95%
Qwen 14B 8K 61.34% 46.33% 31.58% 18.56% 93.00% 69.96% 53.46%
Yi 34B 200K 66.39% 55.13% 38.82% 42.27% 95.00% 83.75% 63.56%
Alfred 40B 8K 52.38% 48.39% 25.00% 27.84% 87.33% 67.14% 51.35%
Beluga 70B 4K 61.90% 46.33% 3.28% 21.65% 81.00% 75.27% 48.24%
Tulu2 70B 8K 64.99% 53.37% 41.45% 34.02% 95.67% 82.33% 61.97%
Jina Embedding Retrieval
Chatglm 6B 32K 52.94% 44.57% 27.63% 23.71% 83.33% 71.38% 50.60%
Mistral 7B 32K 54.90% 43.99% 32.24% 25.75% 79.00% 76.33% 52.04%
Zephyr 7B 32K 52.66% 46.33% 30.92% 23.71% 91.00% 78.09% 53.79%
Longchat 13B 16K 42.58% 33.43% 22.37% 13.40% 57.67% 57.24% 37.78%
Qwen 14B 8K 60.50% 46.63% 44.08% 24.74% 94.33% 78.45% 58.12%
Yi 34B 200K 66.67% 54.25% 45.39% 38.14% 95.00% 83.39% 63.81%
Alfred 40B 8K 50.42% 44.28% 27.63% 25.77% 88.33% 71.02% 51.24%
Beluga 70B 4K 59.94% 49.85% 23.68% 27.84% 96.00% 77.03% 55.72%
Tulu2 70B 8K 64.99% 54.25% 38.82% 31.96% 95.00% 84.10% 61.52%
Table 2: The evaluation results of models on Marathon benchmark. C&R refers to Comprehension and Reasoning task; MIR refers to Multiple Information Retrieval task; TR refers to Timeline Reorder task; Com. refers to Computation task; PR refers to Passage Retrieval task; SDQA refers to Short Dependency Question Answering task; Avg. denotes the average accuracy across all tasks.
4 Experiments
4.1 Setup
Models.
In this evaluation, we selected models with varying parameter sizes, ranging from 7B to 70B, and different context window sizes, ranging from 8k to 200k. The models evaluated in this study include ChatGLM3-6B-32K (Zeng et al., 2022; Du et al., 2022), Mistral-7B-Instruct-v0.1 (Jiang et al., 2023a), Zephyr-7B-$\beta$ (Tunstall et al., 2023), Longchat-13B-16K (Li et al., 2023a), Qwen-14B-Chat (Bai et al., 2023a), Yi-34B (01.AI, 2023), Alfred-40B-1023 (Hallström et al., 2023), StableBeluga-2-70B (Mahan et al., 2023), Tulu-2-DPO-70B (Ivison et al., 2023), ChatGPT-1106 (OpenAI, 2023a), and GPT-4-1106-preview (OpenAI, 2023b).
Methods.
In this evaluation, we first assessed the inherent ability of various models to comprehend long contexts. Then, we evaluated the current mainstream methods for handling long contexts: Compression and RAG. Specifically, for the Compression method, we assessed LongLLMLingua (Jiang et al., 2023b), while for the RAG method, we evaluated two retrieval approaches using OpenAI Embedding and Jina Embedding respectively.
Type Chatglm Mistral Zephyr Longchat Qwen Yi Alfred Beluga Tulu2
Vanilla
JSON 69.48% 77.22% 84.51% 92.29% 62.29% 38.95% 13.27% 22.48% 30.72%
JSON-like 30.52% 21.18% 6.86% 3.99% 0.72% 28.56% 81.96% 0.33% 46.47%
Plain Text 0.00% 6.60% 8.63% 3.73% 36.99% 32.48% 4.77% 71.19% 22.81%
LongLLMLingua Compression
JSON 94.58% 68.10% 63.20% 93.92% 90.92% 48.10% 13.66% 29.97% 35.45%
JSON-like 5.42% 22.68% 9.15% 2.81% 0.06% 26.60% 85.95% 0.59% 43.46%
Plain Text 0.00% 9.22% 27.65% 32.68% 9.02% 25.29% 0.39% 69.54% 19.08%
OpenAI Embedding Retrieval
JSON 52.88% 84.31% 21.83% 31.11% 65.62% 67.71% 16.27% 6.27% 98.43%
JSON-like 42.42% 6.67% 69.87% 23.73% 16.93% 32.16% 83.73% 0.00% 1.11%
Plain Text 4.71% 9.02% 8.30% 45.16% 17.45% 0.13% 0.00% 93.73% 0.46%
Jina Embedding Retrieval
JSON 86.34% 83.14% 17.91% 32.75% 63.33% 65.69% 0.00% 8.43% 97.19%
JSON-like 9.15% 6.21% 73.86% 21.70% 17.91% 34.18% 100.00% 0.007% 2.16%
Plain Text 4.51% 10.65% 8.24% 45.56% 18.76% 0.13% 0.00% 91.50% 0.65%
Table 3: The evaluation results of large language models on the Marathon benchmark for instruction following.
4.2 Implementation Details
Prompt.
We used the same prompt template to ask questions for all models, and required the answers to be returned in JSON format. The specific prompt format can be seen in Figure 9.
LongLLMLingua.
For LongLLMLingua, we set the compression rate to 0.5, the dynamic context compression ratio to 0.4, We also sort the compressed contexts based on their importance.
Embedding Retreival.
For Embedding Retrieval, we utilize the ServiceContext and VectorStoreIndex of the Llama-Index (Liu, 2022). We employ various models as LLMs (Language Models), testing the OpenAI Embedding model and the Jina Embedding model as Embedding Models respectively. The default parameter settings are retained, with a chunk size of 1024 and a top-k value of 2. As for Jina Embedding, we set the pooling method to "mean" to align with Jina’s encode implementation.
Hardware.
All experiments in this evaluation were conducted on a server with 4*A100 80GB.
4.3 Results
4.3.1 Main results
Table 2 presents the performance of different models with different optimization methods on various tasks in the evaluation. For a more intuitive comparison of the results, one can refer to Figure 3, Figure 4, Figure 5, Figure 6, Figure 7, and Figure 8 in appendix. It can be observed that both OpenAI Embedding Retrieval and Jina Embedding Retrieval outperform LongLLMLingua.
In addition, all models have lower accuracy in the Timeline Reorder task and Computation task compared to other tasks. The adoption of the LongLLMLingua method did not lead to improvement, and the improvement achieved by the optimization method RAG was also limited.
Vanilla.
Among the Vanilla methods, the highest accuracy is achieved by Yi-34B, which has 34B parameters but achieves an accuracy of 55.91%. Followed by ChatGLM3-6B-32K, which achieves an accuracy of 55.05% despite having only 6B parameters. Next is Beluga-70B, which achieves an accuracy of 49.51% despite having a context window of only 4K. The average accuracy of the remaining models does not differ significantly, and none of them exceeds 40%.
LongLLMLingua.
Compared to the Vanilla method, LongLLMLingua only improved the accuracy of Qwen by 4.85%, Alfred by 1.51%, Beluga by 3.08% and Tulu2 by 8.64%. However, it reduced the accuracy of ChatGLM3 by 7.14%, Mistral by 2.8%, Zephyr by 7.74%, Longchat by 0.26%, and Yi by 7.25%.
OpenAI Embedding Retrieval.
Compared to the Vanilla method, OpenAI Embedding Retrieval improves the accuracy of Mistral by 10.37%, Zephyr by 11.66%, Qwen by 14.19%, Yi by 7.65%, Alfred by 14.05% and Tulu2 by 24.05%. However, it results in a decrease in the accuracy of ChatGLM3 by 4.06%, Longchat by 5.92% and Beluga by 1.27%.
Jina Embedding Retrieval.
Compared to the Vanilla method, Jina Embedding Retrieval has improved the accuracy of most models. Specifically, Mistral’s accuracy has improved by 12.23%, Zephyr’s accuracy has improved by 15.82%, Longchat’s accuracy has improved by 1.91%, Qwen’s accuracy has improved by 18.15%, Yi’s accuracy has improved by 7.9%, Alfred’s accuracy has improved by 13.93%, Beluga’s accuracy has improved by 6.21%, and Tulu2’s accuracy has improved by 23.60%.
4.3.2 Instruction Following Capability
During the evaluation process, we found that many models had poor instruction-following capabilities. We specified in the prompt that the large language model should respond in JSON format and provided an example of the desired response. However, sometimes the model still responded in other formats or attempted to respond in JSON format but with incomplete results. We conducted a statistical analysis of the evaluation results of various models to estimate their instruction following capabilities, the results are displayed in table 3.
JSON represents the correct generation of a JSON format response by the large language model. JSON-like refers to the attempt by the large language model to generate a response in JSON format but, for various reasons (such as exceeding length limits and being truncated, or having formatting errors), it fails to generate a correct JSON format response. Plain Text indicates that the large language model uses a different format for the response.
From Table 3, we can see that even Yi, which achieved the highest answer accuracy rate, only had an instruction-following capability of 38.95%. Beluga, on the other hand, had an instruction-following capability of only 22.48%. However, Longchat, despite having a low question-answering accuracy rate, exhibited strong instruction-following capabilities with a score of 92.29%, second only to ChatGPT’s 99.61%. Mistral and Zephyr also showed instruction-following capabilities of 77.22% and 84.51%, respectively.
5 Discussion
JSON Format.
At the recent OpenAI Developer Day [^3] https://devday.openai.com, OpenAI announced two major updates for GPT-4. One is parallel function invocation and the other is the ability to specify response format, requiring GPT-4 to return responses in JSON format. These two features are fundamental for building AGI (Artificial General Intelligence). With parallel function invocation, large language models can simultaneously call multiple utility functions to complete user tasks. The ability to respond in JSON format ensures the smooth inclusion of parameters and result returns for function invocation.
Document Context.
After the OpenAI Developer Day, GPT-4 acquired a new feature called Knowledge Retrieval, which allows users to upload their own files and retrieve relevant information from these files to respond to user queries, which is an important application scenario for RAG.
In the future, more and more large language models will implement similar functionalities. The evaluation format for long context question answering tasks will also change. Instead of directly inputting the long context in the prompt, the benchmarking work for long-context question-answering should be aligned with the direction of answering user questions based on the files provided by the user.
Online Benchmarks.
With the continuous development of large language models, the methods and benchmarks for evaluating them also need to be changed accordingly. In the past, when evaluating large language models, we relied on a research team or individual to propose a static benchmark, and then conducted evaluations based on this benchmark. However, the data in static benchmarks would leak over time and be used for model training, resulting in less accurate evaluation results. Moreover, it is inefficient for a research team or individual to build benchmarks, and it is also difficult for researchers to continuously find new evaluation data.
The power of an individual is limited, but the power of a collective is infinite. We call for the construction of Online Benchmarks, which will gather the strength of the entire open-source community to collectively maintain a continuously updated benchmark. This benchmark will have regularly updated evaluation tasks and periodically replaced evaluation formats to cope with the rapid development of large language models.
6 Conclusion
In this experiment, we compared six open source large language models with different parameter sizes and different context windows, as well as OpenAI’s ChatGPT and GPT-4. We also compared the optimization methods based on Compression, LongLLMLingua, and the optimization methods based on RAG, OpenAI Embedding Retrieval, and Jina Embedding Retrieval.
The experimental results show that the method based on RAG is more effective than the method based on Prompt Compression, and Jina Embedding performs better than OpenAI Embedding. Furthermore, the instruction-following ability of open-source large language models is generally poor, despite their ability to achieve high accuracy in question-answering tasks.